OpenStudy (haleyelizabeth2017):

Find cos x if (sin^(2)x -1)/cos x = -1

OpenStudy (anonymous):

\[\frac{ \sin ^2x-1 }{ \cos x }=-1,\frac{ -\left( 1-\sin ^2x \right) }{ \cos x }=-1\] \[\frac{ -\cos ^2x }{ \cos x }=-1\] \[\cos x=1\]

OpenStudy (haleyelizabeth2017):

okay, thank you

OpenStudy (anonymous):

yw

OpenStudy (haleyelizabeth2017):

Sorry for the many questions...LOL

OpenStudy (anonymous):

no problem.